# Causality for Complex Continuous-time Functional Longitudinal Studies with Dynamic Treatment Regimes

### v1

@@ Abstract
Causal inference in longitudinal studies is often hampered by treatment-confounder feedback. Existing methods typically assume discrete time steps or step-like data changes, which we term “regular and irregular functional studies,” limiting their applicability to studies with continuous monitoring data, like intensive care units or continuous glucose monitoring. These studies, which we formally term “functional longitudinal studies,” require new approaches. Moreover, existing methods tailored for “functional longitudinal studies” can only investigate static treatment regimes, which are independent of historical covariates or treatments, leading to either stringent parametric assumptions or strong positivity assumptions. This restriction has limited the range of causal questions these methods can answer and their practicality. We address these limitations by developing a nonparametric framework for functional longitudinal data, accommodating dynamic treatment regimes that depend on historical covariates or treatments, and may or may not depend on the actual treatment administered. To build intuition and explain our approach, we provide a comprehensive review of existing methods for regular and irregular longitudinal studies. We then formally define the potential outcomes and causal effects of interest, develop identification assumptions, and derive g-computation and inverse probability weighting formulas through novel applications of stochastic process and measure theory. Additionally, we compute the efficient influence curve using semiparametric theory. Our framework generalizes existing literature, and achieves double robustness under specific conditions. Finally, to aid interpretation, we provide sufficient and intuitive conditions for our identification assumptions, enhancing the applicability of our methodology to real-world scenarios.
62D20,
62M99,
Dynamic Treatment Regimes,
Causal Inference,
Stochastic Process,
Functional Data,
Semiparametric Theory,
keywords:
[class=MSC]
keywords:
@@
1 Introduction
1.1 Background
Establishing causality is a fundamental goal in scientific research, particularly in fields such as medicine and social sciences. While randomized controlled trials are considered the gold standard for causal inference, they may not always be feasible or ethical. Observational studies, especially those involving complex longitudinal data, pose unique challenges for causal inference due to the presence of time-varying confounders and the need to account for the dynamic nature of treatment and outcome processes.
Functional longitudinal data, characterized by continuous-time processes and high-dimensional measurements, present additional complexities for causal inference. The presence of an infinite number of potential treatment-confounder feedback loops and the lack of a joint density for these processes require novel methodological approaches. Existing methods developed for regular and irregular longitudinal data may not be directly applicable to functional longitudinal data, necessitating the development of a rigorous theoretical framework tailored to this setting.
Traditional methods of adjustment, such as regression models that incorporate a history of treatments and confounders, often fall short in these scenarios. These methods typically assume that confounders are static and do not take into account the dynamic feedback loop between treatment and confounding variables over time. Intervening on past treatment values, a critical aspect in such studies, necessitates accounting for changes in future confounders. These future confounders must then be carefully conditioned upon to correct for confounding. This creates a complex interplay of variables across different time points, complicating the process of establishing clear causal links.
Given these challenges, there is a pressing need for more sophisticated and nuanced methods of causal inference that can effectively handle the intricacies of longitudinal data. Developing these methods is not just an academic exercise; it has profound implications for how we understand and intervene in a range of fields, from healthcare to social policy. As such, the pursuit of advancing causal inference techniques in the context of complex longitudinal studies is both a formidable challenge and an imperative for continued progress in scientific research. In light of these challenges, this paper aims to develop a comprehensive theoretical framework for causal inference in functional longitudinal data under all dynamic treatment regimes, filling an important gap in the current literature.
1.1.1 Longitudinal data formats
To gain a deeper insight into the complexities of functional longitudinal data, it is crucial to first examine the various formats of longitudinal data and their unique features. These datasets can be classified into three primary types. The first type, ”Regular Longitudinal Data (RLD),” involves data collected at set, regular intervals, though this may not accurately reflect the unpredictable nature of visit timings. The second type, ”Irregular Longitudinal Data (ILD),” consists of data where changes in treatment and confounders occur in discrete, finite steps, resembling point processes. This type is common in pharmacoeconomic studies, where patient visits and treatment adjustments happen randomly but within finite periods. The third type, ”Functional Longitudinal Data (FLD),” has emerged with advancements in real-time monitoring technologies in healthcare, characterized by continuous tracking of confounders and treatments. This type of data is common in intensive care and chronic disease management using wearable devices, which results in numerous feedback loops between treatment and confounders, thereby complicating causal inference due to the absence of a joint density for these variables. Figures 1 and 2 provide a visual representation of the distinct characteristics and relationships of these data types.
Figure 1: Possible realizations from three type of longitudinal data. The top left is from a regular longitudinal study with fixed visit times. The top right is from a irregular longitudinal study with irregular and random visit times. The bottom one is from a functional longitudinal study with real-time monitoring data. Reprinted from Ying (2024).

### v2

# Andrew Ying Irvine, CA 92606, USA

@@ Abstract
Establishing causality is a fundamental goal in fields like medicine and social sciences. While randomized controlled trials are the gold standard for causal inference, they are not always feasible or ethical. Observational studies can serve as alternatives but introduce confounding biases, particularly in complex longitudinal data, where treatment-confounder feedback complicates analysis. The challenge increases with Dynamic Treatment Regimes (DTRs), where treatment allocation depends on rich historical patient data. The advent of real-time healthcare monitoring technologies, such as MIMIC-IV and Continuous Glucose Monitoring (CGM), has popularized Functional Longitudinal Data (FLD). However, there is yet no investigate of causal inference for FLD with DTRs. In this paper, we address it by developing a population-level framework for functional longitudinal data, accommodating DTRs. To that end, we define the potential outcomes and causal effects of interest. We then develop identification assumptions, and derive g-computation, inverse probability weighting, and doubly robust formulas through novel applications of stochastic process and measure theory. We further show that our framework is nonparametric and compute the efficient influence curve using semiparametric theory. Last, we illustrate our framework’s potential through Monte Carlo simulations.
1 Introduction
Understanding treatment effects and establishing causality are fundamental goals in fields such as medicine and social sciences. While double-blinded randomized controlled trials are considered the gold standard for causal inference, they may not always be feasible or ethical. Observational studies are sometimes feasible alternative for comparative studies yet introduced additional confounding bias. Complex observational longitudinal data, pose additional challenges for causal inference due to the interplay among time-varying confounders, treatment, and outcome processes, also known as the treatment-confounder feedbacks (Robins, 1998, 1999; Hernán & Robins, 2020). Another complication within longitudinal studies, due to the many possible treatment combinations across time and rich historical information of patient characteristics, treatment process, and prognoses, is the need to investigate causal effects under dynamic treatment regimes, that is, treatment allocation might depend on all historical information.
The advent of real-time monitoring technologies in healthcare has led to the continuous-time measurement of patient data. For example, the Medical Information Mart for Intensive Care IV (MIMIC-IV) (Johnson et al., 2023) is a freely accessible electronic health record (EHR) database that records ICU care data, including physiological measurements, laboratory values, medication administration, and clinical events. Another example is Continuous Glucose Monitoring (CGM) (Rodbard, 2016; Klonoff et al., 2017), an increasingly adopted technology for insulin-requiring patients that provides insights into glycemic fluctuations. CGM offers a real-time, high-resolution stream of data, capturing the intricate fluctuations in interstitial fluid glucose levels every few minutes. These examples illustrate the recent prevalence of functional longitudinal data (FLD).
However, there is no investigation of causal inference for dynamic treatment regimes of longitudinal studies with functional data. This is possibly because of the fact that FLD, characterized by continuous-time processes and high-dimensional measurements, present additional complexities for causal inference due to the presence of an infinite number of potential treatment-confounder feedback loops and the lack of a joint density for these processes. Existing methods developed for classical longitudinal data can not be directly applicable to functional longitudinal data. For less distraction, we dedicate a thorough literature review later in Section 6. Therefore, there is a pressing need for more sophisticated and nuanced methods of causal inference that can effectively handle the intricacies of functional longitudinal data.
We aim to fill this gap partially by developing a comprehensive population-level framework for causal inference in functional longitudinal data under dynamic treatment regimes.
To that end, we proceed with the following steps routinely adopted by classical causal inference frameworks for non-functional longitudinal data:
  1. 1.
    We generalize the potential outcome framework (Neyman, 1923; Rubin, 1974; Holland, 1986) and define the causal parameter of interest as a functional of measures on unobserved potential outcomes, for FLD under general DTRs. Here by “general” we mean the DTRs can be either deterministic or stochastic, and be prespecified or allowed to depend on the actual observed TRs;
  2. 2.
    We propose a set of identification assumptions tailored to the FLD setting, to identify the causal parameter of interest defined earlier. Identification methods include a g-computation formula (Greenland & Robins, 1986; Robins, 2000; Bang & Robins, 2005), an inverse probability weighting (IPW) formula Rosenbaum & Rubin (1983); Hernán et al. (2000), and a doubly robust formula (Robins et al., 1994, 1995; Rotnitzky et al., 1998; Kang & Schafer, 2007). These formulas are functionals of measures on observed data;
  3. 3.
    We deepen our understanding the influences of our identification assumptions on the set of observed data measures by showing they do not pose any restrictions and hence being nonparametric;
  4. 4.
    We also provide a first-order understanding of the identification functionals on the the set of observed data measures by deriving the efficient influence curve (EIC) through semiparametric theory (Bickel et al., 1993; van der Vaart, 2000; Tsiatis, 2006; Kosorok, 2008).
While this paper builds a population-level framework, it does not explore estimation or associated inferential results, which are beyond the scope of this study. With that said, our framework does pave the way for future estimation framework, by suggesting feasible estimators, possibly with efficiency. It also suggest ways to numerically approximate the causal parameter of interest, a non-trivial task for causal inference for FLD, which we illustrate through Monte-Carlo simulations. As a byproduct, our framework subsumes many existing identification methods for non-functional longitudinal data as special cases, thereby providing a unified longitudinal data causal inference framework.
The paper is organized as follows. Section 2 prepares the readers with notation and examples. Section 3 the potential outcome framework and define the causal parameter of interest as a functional of measures on unobserved potential outcomes, for FLD under general DTRs. Section 4 presents our causal identification framework for FLD under ADTR, including the key assumptions and the derivation of the g-computation and inverse probability weighting (IPW) formulas. In this Section we also discuss the recursive representations of the g-computation and IPW processes, presents the doubly robust formula. Section 5 proves that our identification assumptions do not impose any restrictions on the set of observed data distributions. Therein we also develop the semiparametric theory for our proposed estimators, deriving the efficient influence function (EIF) and discussing its implications. We conduct Monte Carlo simulations to examine our framework in Section LABEL:sec:simu, which also implies that our framework suggests a numerical approach to approximate the causal quantity of interest. In Section 6 we give a literature review. Finally, Section 7 concludes the paper with a summary of our main contributions and a discussion of future research directions. In the appendices, Our identification assumptions are based on the so-called “coarsened at random assumption,” which is stronger than the classical “sequential randomization assumption (SRA).” Section A provides an alternative weaker but less interpretable set of identification assumptions extending SRA, under which the results in our paper continue to hold. For readers who are less familiar with longitudinal data method or who are familiar with classical method but interested in drawing connections, we provide a review of existing methods for causal inference mathematically in classical longitudinal data in the supplementary material, highlighting their limitations when applied to FLD. Therein we also review existing approaches (Ying, 2022) designed for FLD, which yet did not accommodate DTRs. Proofs of our results are also given in the supplementary material.
2 Preliminaries
Throughout the rest of the paper, suppose that there is a longitudinal study during 0 to $\infty$. We first define stochastic processes:
  1. 1.
    $A(t)$ are treatments received at time $t$, which can be multi-dimensional binary, categorical or functional. We write $\bar{A}(t)=\{A(s):0\leq s\leq t\}$ and abbreviate $\bar{A}=\bar{A}(\infty)$. We define $\mathcal{A}$ as the set of all possible values of $\bar{a}$.
  2. 2.
    $L(t)$ are measured confounders at time $t$, which can be multi-dimensional binary, categorical, or functional. We write $\bar{L}(t)=\{L(s):0\leq s\leq t\}$ and abbreviate $\bar{L}=\bar{L}(\infty)$. For convenience, we write $\bar{Y}$, a subset of $\bar{L}$, as the outcomes of interest. We include the outcomes into covariates to reflect the fact that both past outcomes and covariates influence current treatment, which can affect both future outcomes and covariates.
  3. 3.
    To accommodate mortality and censoring, typically encountered in longitudinal studies, we consider $T$ as a time-to-event endpoint, for instance, death, and $C$ as the right censoring time, both of which are terminating states. Define $X=\min(T,C)$ as the censored event time. To ease notation, we include $T$ into $\bar{Y}$ and set $C=\infty$ whenever $T\leq C$.
  4. 4.
  5. 5.
    We introduce the counterfactual covariates $L_{\bar{a}}(t)$ and counterfactual outcomes $Y_{\bar{a}}(t)$, for any $\bar{a}\in\mathcal{A}$, as the covariates and outcomes if the treatment were to set as $\bar{a}$. These counterfactuals are defined for deterministic and static treatment regimes $\bar{a}$ (see Section 6.3) and considered to be defined at baseline. We write $\bar{L}_{\mathcal{A}}=\{\bar{L}_{\bar{a}}\}_{\bar{a}\in\mathcal{A}}$ and $\bar{Y}_{\mathcal{A}}=\{\bar{Y}_{\bar{a}}\}_{\bar{a}\in\mathcal{A}}$.
  6. 6.
    The observed data are $(C,\bar{A},\bar{L})=(C,\bar{A}(X),\bar{L}(X))$ and the uncensored full data are $(C,\bar{A}(T_{\mathcal{A}}),\bar{L}_{\mathcal{A}}(T_{\mathcal{A}}))$.
  7. 7.
  8. 8.
    We use the upper case for random variables and the lower case for their realized values.
We then introduce notation for measures:
  1. 1.
    When there is no confusion, we may write $\operatorname{\mathbb{P}}(\textup{d}\bar{a}\textup{d}\bar{l})$ to represent the distribution on the path space induced by the stochastic processes and the probability measure $\operatorname{\mathbb{P}}$ on the sample space $\Omega$. Note that this is not a density. This notation is well adopted by probabilists (Bhattacharya & Waymire, 2007; Durrett, 2019) and also statisticians (Gill & Robins, 2001).
  2. 2.
  3. 3.
  4. 4.
    Note that conditional probability is only uniquely defined almost surely. Therefore throughout this draft, otherwise stated, for the measure zero subset where the conditional probability is not uniquely defined, we set the conditional probability to be zero.
  5. 5.
  6. 6.
    We define Gâteaux derivative. For any two probability distribution $\operatorname{\mathbb{P}}$ and $\operatorname{\mathbb{P}}^{\prime}$ on $(\Omega,\mathscr{F})$, we write $\frac{\partial\operatorname{\mathbb{P}}^{\prime}}{\partial\operatorname{\mathbb{P}}}$ as the Gâteaux derivative of $\operatorname{\mathbb{P}}^{\prime}$ to the direction by $\operatorname{\mathbb{P}}$, understood in the linear space of all signed measure on $(\Omega,\mathscr{F})$.
Finally, a partition $\Delta_{K}[0,\infty]$ on $[0,\infty]$ is defined as a finite sequence of $K+1$ numbers of the form $0=t_{0}<\cdots<t_{K}=\infty$. The mesh $|\Delta_{K}[0,\infty]|$ of a partition $\Delta_{K}[0,\infty]$ is $\max_{i=0,\cdots,K-1}(t_{j+1}-t_{j})$, representing the maximum gap length of the partition.
3 Causal Parameters of Interest
Here we first rigorously define counterfactual variables for any DTRs. For any partition $\Delta_{K}[0,\infty]$, we define a hypothetical treatment $A_{\mathbb{G}_{\Delta_{K}[0,\infty]}}(t)$ as a random draw from
$\bar{A}_{\mathbb{G}_{\Delta_{K}[0,\infty]}}(t_{0})\sim\mathbb{G}\{\textup{d}a(t_{0})\},$ (1)
we then define $Y_{\mathbb{G},\Delta_{K}[0,\infty]}(t_{0})$ as plugging in $\bar{A}_{\mathbb{G}_{\Delta_{K}[0,\infty]}}(t_{0})$ into $Y_{\bar{a}(t_{0})}(t_{0})$ as
$Y_{\mathbb{G},\Delta_{K}[0,\infty]}(t_{0})=Y_{\bar{A}_{\mathbb{G}_{\Delta_{K}[0,\infty]}}(t_{0})}(t_{0}).$ (2)
Therefore we can iteratively define
$\bar{A}_{\mathbb{G}_{\Delta_{K}[0,\infty]}}(t_{1})\sim\mathbb{G}\{\textup{d}\bar{a}(t_{1})|\bar{A}(t_{0}),\bar{L}(t_{0}),Y_{\mathbb{G},\Delta_{K}[0,\infty]}(t_{0})\},$ (3)
$Y_{\mathbb{G},\Delta_{K}[0,\infty]}(t_{1})=Y_{\bar{A}_{\mathbb{G}_{\Delta_{K}[0,\infty]}}(t_{1})}(t_{1}).$ (4)
We iterate until infinity, so we have $\bar{Y}_{\mathbb{G},\Delta_{K}[0,\infty]}$ defined. Now, we can let $|\Delta_{K}[0,\infty]|\to 0$ and define
Definition 1 (Counterfactuals under DTRs).
When the distributions $Y_{\mathbb{G},\Delta_{K}[0,\infty]}$ converge to the same limit in the total variation sense, we define a realization of the limit distribution as $\bar{Y}_{\mathbb{G}}$.
The parameter of interest is defined as
$\operatorname{\mathbb{E}}\{\nu(\bar{Y}_{\mathbb{G}})\},$ (5)
where $\nu$ is a user-specified bounded continuous function. This estimand includes those considering the marginal mean of outcomes under a STR in Ying (2024).
Now we investigate when $\bar{Y}_{\mathbb{G}}$ is well-defined. In fact, when
  1. 1.
@@ 3 Causal Parameters of Interest
  2. 2.
    or $\mathbb{G}_{2}$ is discrete.
Like for Riemannian integral, condition 1 or 2 serves as the corresponding “absolute contuity” or “discrete” for a function to be “Riemannian integrable”. Condition 1 typically holds for continuous-path stochastic process like Wiener measure while condition 2 holds for point measure like Poisson process.
Under either conditions listed above, the distributions of $\bar{Y}_{\mathbb{G},\Delta_{K}[0,\infty]}$ converge in the total variation sense to the same point whenever $|\Delta_{K}[0,\infty]|\to 0$.
  • Though the data are allowed to be functional and the underlying data generating mechanism can have uncountably infinite number of treatment-confounder feedbacks, a finite-dimensional distribution intervention can still be considered, for example, intervening dosage of certain drug hourly or daily;
  • Likewise, a marked point process measure represents intervening both dosage and frequency of usage for certain drugs;
4 Identification
4.1 Identification assumptions
We define
Our first assumption is the generalized “full conditional exchangeability” assumption in causal inference (Robins & Hernan, 2008), which generalized “coarsening at random” assumption in missing data literature (Heitjan & Rubin, 1991).
Assumption 1 (Full sequential randomization).
[…]
This assumption requires that the treatment assignment mechanism at any given time point is independent of all future potential outcomes and covariates, given the observed history up to that time point. To assist Assumption 1, we assume:
Assumption 2 (Full consistency).
$\bar{L}=\bar{L}_{\bar{A}}.$ (12)
It states that, the observed covariate value under any TR is identically distributed with the counterfactual covariate value under the same TR.
We also have the positivity Assumption:
Assumption 3 (Full positivity).
almost surely.
Furthermore, we need an assumption over the censoring mechanism to eliminate the censoring bias. We consider the conditionally independent censoring assumption in Van der Laan & Robins (2003); Rotnitzky & Robins (2005); Tsiatis (2006); Ying (2022). Define the full data censoring time hazard as
$\lambda_{C}\{t|T,\overline{A}(T),\overline{L}(T)\}=\frac{\partial}{\partial\textup{d}t}\operatorname{\mathbb{P}}\{C\leq t+\textup{d}t|C>t,T,\overline{A}(T),\overline{L}(T)\}/\textup{d}t.$ (14)
Assumption 4 (Conditional independent censoring).
The censoring mechanism is said to be conditionally independent if
$\lambda_{C}\{t|T,\overline{A}(T),\overline{L}(T)\}=\lambda_{C}\{t|\mathscr{F}_{t}\}\mathbbm{1}(T>t).$ (15)
4.2 Identification formulas
Definition 2 (G-computation process).
Under Assumptions 1 and 3, define
Theorem 1 (G-computation formula).
Under Assumptions A.1 - 4, the marginal mean of transformed potential outcomes under a user-specified regime $\mathbb{G}$ (5) is identified via a g-computation formula as
Definition 3 (Inverse probability weighting process).
Under Assumptions 1 and 3, define
Theorem 2 (Inverse probability weighting formula).
Under Assumptions A.1 - 4, the marginal mean of transformed potential outcomes under a user-specified regime $\mathbb{G}$ (5) is identified via an inverse probability weighting formula as
We also define $\Xi_{\text{DR},\mathbb{G}}(H,Q)$ as the $L^{2}$ weak limit of $\Xi_{\text{DR},\mathbb{G},\Delta_{K}[0,\infty]}(H,Q)$ whenever it exists.
Proposition 2 (Double robustness).
Under Assumptions A.1 - 4, and .5, and assuming
This theorem presents a doubly robust identification formula for the marginal mean of transformed potential outcomes under a DTR. The doubly robust formula combines the g-computation and IPW processes in a way that provides robustness against model misspecification. Specifically, the theorem shows that the doubly robust formula remains valid if either the g-computation process or the IPW process is correctly specified, even if the other is misspecified. This property enhances the reliability of causal effect estimates in practice.
5 Understanding of Probability Set and Functionals
In this section, we investigate the first order behavior of our identification functional around an observed data distribution $\operatorname{\mathbb{P}}$. To that end, we leverage semiparametric theory. Semiparametric theory, as discussed in key works (Bickel et al., 1993; van der Vaart, 2000; Tsiatis, 2006; Kosorok, 2008; Kennedy, 2017), primarily deals with the local approximation of both the probability measure and its associated functional (or estimand). This theory emerges from applying principles of differential geometry to the concept of a probability measure “manifold.” It is important to distinguish this from semiparametric models, which describe probability distributions using both finite- and infinite-dimensional parameters. In the context of a differentiable manifold with a differentiable real-valued function (estimand) defined on it, the differential is understood as a linear approximation to the function at points along the tangent vectors. These tangent vectors represent an equivalence class of differentiable curves, defined through the equivalence relation of first-order contact between the curves. To calculate the differential, one can use any direction, or equivalently, any tangent vector, to determine the directional derivative. In the case of a probability measure “manifold,” these concepts take on explicit representations. Semiparametric theory characterizes the local neighborhood of a probability measure $\operatorname{\mathbb{P}}$ using the Hellinger distance. Consequently, for any smooth curve, also known as a parametric submodel, that passes through $\operatorname{\mathbb{P}}$, one can derive a corresponding tangent vector, or score, by employing Hadamard derivatives in conjunction with the Hellinger distance. The tangent space is defined as the completion of the linear spans of all tangent vectors, which forms a subspace of $L_{0}^{2}(\operatorname{\mathbb{P}})$. The directional derivative, acting as a continuous linear functional on the tangent vectors, possesses Riesz representers known as influence curves. The projection of these influence curves onto the tangent space is both unique and termed the efficient influence curve (EIC).
Understanding the differential representation of the estimand is crucial in practice. This is because the differential form has Riesz representers, which often suggests viable estimators in numerous scenarios that can substantially mitigate first-order bias. This aspect of semiparametric theory, due to its standalone significance and historical establishment in the literature for RLD and ILD (Rytgaard et al., 2022), underscores the need for a comprehensive semiparametric framework tailored for FLD. Historically, the development of semiparametric theory in the realm of FLD has been notably sparse, possibly attributed to the inherent complexities and challenges in managing FLD without a defined density function. Recognizing this gap, our work endeavors to pioneer in this field by formulating a generalized semiparametric theory for FLD. This advancement not only bridges a significant gap in the existing literature but also opens new avenues for enhanced data analysis and interpretation in studies involving FLD. The novelty of our approach lies in its ability to adapt and extend established principles of semiparametric theory to the unique characteristics and requirements of FLD, thereby offering a more robust and versatile analytical framework that is poised to transform the handling and understanding of such data.
Theorem 3.
For any measure $\operatorname{\mathbb{P}}$ over observed data $(C,\bar{A},\bar{L})$ satisfying Assumption A.7, there exists a measure $\operatorname{\mathbb{P}}^{F}$ over $(C,\bar{A},\bar{L},\bar{L}_{\mathcal{A}})$ satisfying Assumptions A.1 - 4 and inducing $\operatorname{\mathbb{P}}$ over $(C,\bar{A},\bar{L})$.
We show that the tangent space is full, as the space of all $L_{0}^{2}(\operatorname{\mathbb{P}})$ functions. This the “next best” result to full nonparametric property, that is, within the neighborhood of $\operatorname{\mathbb{P}}$ satisfying Assumptions A.1 - 4, the other distributions satisfying Assumptions A.1 - 4 are so rich that one cannot distinguish them with those who do no satisfy in an $L_{0}^{2}$ manner. This helps to establish the limiting IC of any ADTRs $\mathbb{G}$ as $|\Delta_{K}[0,\infty]|\to 0$, is the EIC, that is, the unique IC with shortest $L^{2}$ length and lies on the tangent space. A full tangent space trivially implies that any IC is EIC.
Theorem 4.
Under Assumptions A.1 - 4, the tangent space at $\operatorname{\mathbb{P}}$ is $L_{0}^{2}(\operatorname{\mathbb{P}})$.
Finally we arrive at the EIC-based identification formula. Define
$\Xi_{\mathbb{G}_{\Delta_{K}[0,\infty]}}$
Theorem 5 (Efficient Influence Curve).
Under Assumptions A.2 and A.3, at the law when the derivative and the limit below can interchange,
$\Xi_{\mathbb{G}}-\operatorname{\mathbb{E}}_{\mathbb{G}}\{\nu(\bar{Y})\},$ (25)
where $\Xi_{\mathbb{G}}(H,Q)$ is the $L^{2}$ weak limit of $\Xi_{\mathbb{G}_{\Delta_{K}[0,\infty]}}(H,Q)$ as $|\Delta_{K}[0,\infty]|\to 0$. Furthermore under Assumptions A.1 - 4, the EIC-based formula is
6 Related Work
6.1 Causal Inference for Non-Functional Longitudinal Data
Non-functional longitudinal data can be classified into two primary types:
  • Regular Longitudinal Data (RLD): RLD involves data collected at set, regular intervals, though this may not accurately reflect the unpredictable nature of visit timings.
  • Irregular Longitudinal Data (ILD): ILD consists of data where changes in treatment and confounders occur in discrete, finite steps, resembling point processes. This type is common in pharmacoeconomic studies, where patient visits and treatment adjustments happen randomly but within finite periods.
Figures 1 provides a visual representation of relationships of different longitudinal data types.